# A conjecture on descents, inversions and the weak order on Coxeter groups

### v2

@@ Abstract.
We state the conjecture that the number of right descents of $w$ is the sum of the number of right descents of the elements of $\mathcal{P}$ and prove that this conjecture holds in the cases of symmetric groups (type $A$) and hyperoctahedral groups (type $B$).
@@ 1. Introduction
[…]
@@
  • A $2$-partition of $w$ is called a bipartition of $w$.
  • A partition $\mathcal{P}$ of $w\in W$ is called a proper partition of $w$ if $e\notin\mathcal{P}$; $\{e,w\}$ is a bipartition of $w$ that is not proper.
@@
Let $\{u,v\}$ be a bipartition of $w\in W$, then $d_{R}(w)=d_{R}(u)+d_{R}(v)$.
This conjecture holds for dihedral groups, see Example 1.2 below and for universal Coxeter systems (Proposition 2.3).
Conjecture 1 is a generalization of a conjecture formulated by Ressayre in the case of finite Weyl groups, which the first author was made aware of in 2014. Let $W$ be a finite Weyl group with longest element $w_{\circ}$ and lwt $u_{1},u_{2},u_{3}\in W$, then:
$T(w_{\circ})=T(u_{1})\sqcup T(u_{2})\sqcup T(u_{3})\implies\operatorname{rk}(W)=d_{R}(u_{1})+d_{R}(u_{2})+d_{R}(u_{3})\qquad(\triangle).$
In 2023, Francone and Ressayre [10, Theorem 1] showed that the structure coefficients of the Belkale-Kumar product are either 00 or $1$ for any finite Weyl group, answering by the positive [5, Claim 1]. Their theorem implies ()($\triangle$)( △ ) as a corollary [10, Corollary 5]. Their proof relies on an algebraic geometry argument, based on the fact that complete flag varieties are simply connected, to reduce the problem to a case-by-case computation on crystallographic root systems spanning over twenty pages. It turns out that Conjecture 1 is equivalent to statement ()($\triangle$)( △ ), see Proposition 2.7. So Conjecture 1 is true for finite Weyl groups as a consequence of [10, Theorem 1]. However, no direct [^1] A proof involving solely argument from the theory of Coxeter groups. proof is known and, still, we have made numerous computations with Sagemath [11] to check the validity of Conjecture 1 in infinite Coxeter groups. In this article we provide a direct proof of Conjecture 1, hence of statement ()($\triangle$)( △ ), for type $A$ and type $B$.
@@
$s$$t$$m$
If $m\not=\infty$, only the longest element $w_{\circ}=sts\cdots=tst\cdots$ ($m=\ell(w_{\circ})$ letters) admits proper bipartitions. Indeed, if $w\in W$ admits a proper partition $\mathcal{P}$ with $|\mathcal{P}|>1$, then each for $u\in\mathcal{P}$, $T(u)$ contains at least an element of $S$ (a left descent), since $e\not\in\mathcal{P}$. Because the union of the inversion sets of elements in a partition is disjoint, $|\mathcal{P}|=2$ and $s,t\in T(u)\sqcup T(v)=T(w)$. So $w=w_{\circ}$. All the proper bipartitions of $w_{\circ}$ are therefore of the form $\{u=st\cdots,v=ts\cdots\}$ with $\ell(u)=i$, $\ell(v)=m-i$ and $1\leq i<m$. Therefore we have $d_{R}(w_{\circ})=2=1+1=d_{R}(u)+d_{R}(v)$. In particular, the set of partition-irreducible elements is $W\setminus\{w_{\circ}\}$ in this case.
The same argument that above shows that, if $m=\infty$, then all the elements of $W$ are partition-irreducible.
@@
$T(w)=\{\tau_{1},\tau_{(13)}\}\sqcup\{\tau_{(14)},\tau_{(24)},\tau_{(34)}\}=T(\tau_{1}\tau_{2})\sqcup T(\tau_{3}\tau_{2}\tau_{1})=T(u)\sqcup T(v).$
@@
$T(w^{\prime})=\{\tau_{1}\}\sqcup\{\tau_{(14)},\tau_{(24)},\tau_{(34)}\}=T(\tau_{1})\sqcup T(v).$
@@
If $\{u,v\}$ s a bipartition of $w$, $D_{R}(w)\not=D_{R}(u)\sqcup D_{R}(v)$ in general, as observed in Example 1.3 above.
In this article we provide a direct proof of Conjecture 1 for Coxeter systems of type $A$ (symmetric groups) and $B$ (hyperoctahedral groups). For $n\in\mathbb{N}^{*}$, we denote by $\mathcal{S}_{n}$ the symmetric group of permutations of $\{1,\dots,n\}$ and by $W_{n}$ the hyperoctahedral group of signed permutations of $[\![\bar{n},n]\!]^{*}:=\{\pm 1,\dots,\pm n\}$.
Theorem 1.5.
$d_{R}(w)=d_{R}(u)+d_{R}(v).$
Moreover, if $\{u_{1},\dots,u_{k}\}$ is a $k$-partition of $w$, then $d_{R}(w)=\sum_{i=1}^{k}d_{R}(u_{i})$.
The proof of the first part of Theorem 1.5 is given in §3 (for type $A$) and in §4 (for type $B$). The second part follows from a more general consequence of Conjecture 1, which is proven in §2.
Proposition 1.6.
We have made numerous computations with Sagemath [11] to check the validity of Conjecture 1; see §5 for details. These computations were made using an equivalent description for bipartitions of elements in term of interval in the (right) weak order.
Recall that the (right) weak order is the poset $(W,\leq_{R})$ defined by $u\leq_{R}w$ if and only if $u$ is a prefix of $w$, i.e., a reduced word for $u$ appear as the prefix of a reduced word for $w$. This is equivalent to consider the right Cayley graph of $(W,S)$ with the edges oriented as followed: an edge labeled by $s\in S$ is oriented from $w$ to $ws$ if and only if $s\notin D_{R}(w)$, if and only if $\ell(ws)>\ell(w)$.
@@
Let $w\in W$. A subset $\{u,v\}$ of the interval $[e,w]_{R}$ in $(W,\leq_{R})$ is a diameter of $[e,w]_{R}$ if $d(u,v)=\ell(w)$. In particular, the pair $\{e,w\}$ is a diameter. The interval $[e,w]_{R}$ is called rectangled if it has at least two diameters.
@@
Let $u,v,w\in W$, then $\{u,v\}$ is a bipartition of $w$ if and only if $\{u,v\}$ is a diameter of $[e,w]_{R}$. In particular, the interval $[e,w]_{R}$ is rectangled if and only if $w$ is not partition-irreducible.
The above proposition allows for two useful reformulations of Conjecture 1. For a non-empty interval $[u,v]_{R}$ in the right weak order $(W,\leq_{R})$, we denote by:
  • $\operatorname{atom}([u,v]_{R})$ the number of atoms in $[u,v]_{R}$, i.e., the number of elements in $[u,v]_{R}$ that cover $u$;
  • $\operatorname{coatom}([u,v]_{R})$ the number of coatoms in $[u,v]_{R}$, i.e., the number of elements in $[u,v]_{R}$ that are covered by $v$.
Since $d_{R}(g)=\operatorname{coatom}([e,g]_{R})$ for any $g\in W$, Conjecture 1 is equivalent to the following conjecture.
@@
Let $u,v,w\in W$ such that $\{u,v\}$ is a diameter of $[e,w]_{R}$. Then $\operatorname{coatom}([e,w]_{R})=\operatorname{coatom}([e,u]_{R})+\operatorname{coatom}([e,v]_{R})$.
It is well-known that the left multiplication by $w^{-1}$ is an order-reversing isomorphism from $[e,w]_{R}$ to $[e,w^{-1}]_{R}$, i.e., $u\leq_{R}v$ if and only if $w^{-1}v\leq_{R}w^{-1}u$, for all $u,v\in[e,w]_{R}$. In particular, $\operatorname{coatom}([e,w]_{R})=\operatorname{atom}([e,w^{-1}]_{R}$ and, for any $u\leq_{R}w$, $\operatorname{coatom}([e,u]_{R})=\operatorname{atom}([w^{-1}u,w^{-1}]_{R}$. So Conjecture 2, and hence Conjecture 1, are equivalent to the following conjecture; see Proposition 2.14 for more details.
@@
Let $u,v,w\in W$ such that $\{u,v\}$ is a diameter of $[e,w]_{R}$. Then $\operatorname{atom}([e,w]_{R})=\operatorname{atom}([u,w]_{R})+\operatorname{atom}([v,w]_{R})$.
Together with Theorem 1.5 we obtain therefore the following corollary.
@@
Let $n\in\mathbb{N}^{*}$ and assume that $W=\mathcal{S}_{n}$ or $W=W_{n}$. Then for any $w\in\mathcal{W}$ and $\{u,v\}$ a diameter of $[e,w]_{R}$, we have:
@@
Refer to caption1111333344442222
Figure 1. The rectangled interval $[e,42131]_{R}$ for $(W,S)$ of type $D_{4}$ with Coxeter graph on the upper left side. This rectangled interval has only two diameters: $\{e,42131\}$ and $\{1,4231\}$. The coatoms of $[e,42131]_{R}$ are $4213,4231$, the coatom of $[e,1]_{R}$ are $e$ and the coatoms of $[e,4231]_{R}$ is $423$, so Conjecture 2, and therefore Conjecture 1, hold in this example. The atoms of $[e,42131]_{R}$ are $1,2,4$ (the left descents), the atoms of $[1,42131]_{R}$ are $41,21$ and the atom of $[4231,42131]_{R}$ is $42131$, which illustrates the statement of Conjecture 3.
@@ Acknowledgements
Part of this research was done at the Laboratoire d’algèbre, de combinatoire et d’informatique mathématique (LACIM) at UQAM during the sabbatical of the second author. Therefore, she warmly thanks the IRL CRM-CNRS of Montréal and the INSMI for giving her this opportunity. The first author warmly thanks Riccardo Biagioli, Matthew Dyer, Philippe Nadeau, Piotr Przytycki, Nicolas Ressayre and Paolo Santinelli for instructive discussions.
@@ 2. Partitions of elements in a Coxeter system
In the weak order $(W,\leq_{R})$, it is well-known that, for $u,w\in W$, we have $u\leq_{R}w$ if and only if $T(u)\subseteq T(w)$. Börner showed that $(W,\leq_{R})$ is a complete meet-semilattice. For $X\subseteq W$, we denote the meet of $X$ by $\bigwedge_{R}X$ and, if it exists, the join of $X$ by $\bigvee_{R}X$.
@@ 2.1. First properties of partitions of elements of $W$
(1) is a direct consequences of the definition of partitions of $w$. (2) is a consequence of the definition of partitions of $w$ and of the fact that $|T(g)|=\ell(g)$ for any $g\in W$. (3) is a consequence of the definition of partitions of $w$ and of the fact that $u\leq_{R}w$ if and only if $T(u)\subseteq T(w)$. (4) Assume by contradiction that $\wedge_{R}\mathcal{P}\not=e$. Then there is $s\in S$ such that $s\leq_{R}u$ for all $u\in\mathcal{P}$. So
@@
Recall that $(W,S)$ is a universal Coxeter system if its Coxeter graph is a complete graph whose edges are all indexed by $\infty$. The following proposition states that universal Coxeter systems satisfies Conjecture 1, which is obviously true if there is only partition-irreducible elements in $W$.

### v3

@@ Abstract.
Moreover, we state and discuss the conjecture that the number of right descents of $w$ is the sum of the number of right descents of the elements of $\mathcal{P}$. In particular: (1) we state an equivalent conjecture in term of the number of atoms in the interval $e,w]_{R}$ of the right weak order, with a word-metric flavour; (2) we give a direct proof that this conjecture holds in the cases of symmetric groups (type $A$) and hyperoctahedral groups (type $B$).
@@ 1. Introduction
$T(w)=\bigsqcup_{u\in\mathcal{P}}T(u)\quad\textrm{(disjoint union)}.$
@@
  • A partition $\mathcal{P}$ of $w\in W$ is called a proper partition of $w$ if $e\notin\mathcal{P}$; $\{e,w\}$ is a $2$-partition of $w$ that is not proper.
  • A proper $2$-partition of $w$ is called a bipartition of $w$.
@@
Let $(W,S)$ be a Coxeter system. If $\{u,v\}$ is a bipartition of $w\in W$, then $d_{R}(w)=d_{R}(u)+d_{R}(v)$.
Conjecture 1 is a generalization of a conjecture formulated by Ressayre in the case of finite Weyl groups, which the first author was made aware of in 2014; see §1.2 for more details on Ressayre’s original conjecture.
The validity of Conjecture 1 guarantees such an equality of sum of right descent sets also for arbitrary partitions of an element of $W$; the following proposition is proven in §2.
Proposition 1.2.
As far as we know, Conjecture 1 holds for dihedral groups (Example 1.3 below), for universal Coxeter systems, since all elements in that case are partition-irreducible (Proposition 2.3), as well as for finite Weyl group, as a consequence of [10]. However, in the case of finite Weyl group, no direct [^1] A proof involving solely argument from the theory of Coxeter groups. proof of Conjecture 1 is known, see the discussion in §1.2.
In this article we provide a direct proof of Conjecture 1 for type $A$ in §3 and for type $B$ in §4, which we believe have a some potential to be generalized to arbitrary Coxeter systems (see Remark 3.2). We have also made numerous computations with Sagemath [11] to check the validity of Conjecture 1 in infinite Coxeter groups, see §5.
@@
ssttmm
If $m\not$$\infty$, only the longest element $w_{\circ}=sts\cdots=tst\cdots$ ($m=\ell(w_{\circ})$ letters) admits proper bipartitions. Indeed, if $w\in W$ admits a proper partition $\mathcal{P}$ with $|\mathcal{P}|>1$, then each for $u\in\mathcal{P}$, $T(u)$ contains at least an element of $S$ (a left descent), since $e\not\in\mathcal{P}$. Because the union of the inversion sets of elements in a partition is disjoint, $|\mathcal{P}|=2$ and $s,t\in T(u)\sqcup T(v)=T(w)$. So $w=w_{\circ}$. All the proper bipartitions of $w_{\circ}$ are therefore of the form $\{u=st\cdots,v=ts\cdots\}$ with $\ell(u)=i$, $\ell(v)=m-i$ and $1\leq i<m$. Therefore we have $d_{R}(w_{\circ})=2=1+1=d_{R}(u)+d_{R}(v)$. In particular, the set of partition-irreducible elements is $W\setminus\{w_{\circ}\}$ in this case.
The same argument as above shows that, if $m=\infty$, then all elements of $W$ are partition-irreducible.
@@
$T(w)=\{\tau_{1},\tau_{(13)}\}\sqcup\{\tau_{(14)},\tau_{(24)},\tau_{3}\}=T(\tau_{1}\tau_{2})\sqcup T(\tau_{3}\tau_{2}\tau_{1})=T(u)\sqcup T(v).$
@@
$T(w^{\prime})=\{\tau_{1}\}\sqcup\{\tau_{(14)},\tau_{(24)},\tau_{3}\}=T(\tau_{1})\sqcup T(v).$
@@
If $\{u,v\}$ is a bipartition of $w$, $D_{R}(w)\not=D_{R}(u)\sqcup D_{R}(v)$ in general, as observed in Example 1.4 above.
1.1. Conjecture 1 and the right weak order
The computations to check the validity of Conjecture 1 with Sagemath [11], for which we present some enumerative results in §5, were made using an equivalent description for bipartitions of elements in intervals in the (right) weak order, endowed with the word-metric. Recall that the (right) weak order is the poset $(W,\leq_{R})$ defined by $u\leq_{R}w$ if and only if $u$ is a prefix of $w$, i.e., a reduced word for $u$ appear as the prefix of a reduced word for $w$. This is equivalent to consider the right Cayley graph of $(W,S)$ with the edges oriented as followed: an edge labeled by $s\in S$ is oriented from $w$ to $ws$ if and only if $s\notin D_{R}(w)$, if and only if $\ell(ws)>\ell(w)$.
@@ 1.1. Conjecture 1 and the right weak order
Let $w\in W$. A subset $\{u,v\}$ of the interval $e,w]_{R}$ in $(W,\leq_{R})$ is a diameter of $e,w]_{R}$ if $d(u,v)=\ell(w)$. In particular, the pair $\{e,w\}$ is a diameter. The interval $e,w]_{R}$ is called rectangled if it has at least two diameters.
@@
Let $u,v,w\in W$, then $\{u,v\}$ is a bipartition of $w$ if and only if $\{u,v\}$ is a diameter of $e,w]_{R}$. In particular, the interval $e,w]_{R}$ is rectangled if and only if $w$ is not partition-irreducible.
The above proposition allows for two useful reformulations of Conjecture 1. For a non-empty interval $u,v]_{R}$ in the right weak order $(W,\leq_{R})$, we denote by:
  • $\operatorname{atom}([u,v]_{R})$ the number of atoms in $u,v]_{R}$, i.e., the number of elements in $u,v]_{R}$ that cover $u$;
  • $\operatorname{coatom}([u,v]_{R})$ the number of coatoms in $u,v]_{R}$, i.e., the number of elements in $u,v]_{R}$ that are covered by $v$.
@@
Let $u,v,w\in W$ such that $\{u,v\}$ is a diameter of $e,w]_{R}$. Then $\operatorname{coatom}([e,w]_{R})=\operatorname{coatom}([e,u]_{R})+\operatorname{coatom}([e,v]_{R})$.
@@
Let $u,v,w\in W$ such that $\{u,v\}$ is a diameter of $e,w]_{R}$. Then $\operatorname{atom}([e,w]_{R})=\operatorname{atom}([u,w]_{R})+\operatorname{atom}([v,w]_{R})$.
Since $d_{R}(g)=\operatorname{coatom}([e,g]_{R})$ for any $g\in W$, Conjecture 1 is equivalent to Conjecture 2, by Proposition 1.7. However, Conjecture 3 is a switch of perspective: it states if true that the number of left descents of $w$ is equal to the number of right ascents of $u$ plus the number of right ascents of $v$ such that the geodesics they start remains in the interval $e,w]_{R}$. The following theorem is proven in §2.5.
Theorem 1.8.
Conjecture 1, Conjecture 2 and Conjecture 3 are equivalent.
Since Conjecture 1 holds for finite Weyl groups, in particular for symmetric and hyperoctahedral groups, we obtain the following corollary, for which it would be interesting to have a proof involving only weak order technics.
@@
Assume that $W$ is a finite Weyl group. Then for any $w\in W$ and $\{u,v\}$ a diameter of $e,w]_{R}$, we have:
@@
Refer to caption11334422
Figure 1. The rectangled interval $e,42131]_{R}$ for $(W,S)$ of type $D_{4}$ with Coxeter graph on the upper left side. This rectangled interval has only two diameters: $\{e,42131\}$ and $\{1,4231\}$. The coatoms of $e,42131]_{R}$ are $4213,4231$, the coatom of $e,1]_{R}$ is $e$ and the coatoms of $e,4231]_{R}$ is $423$, so Conjecture 2, and therefore Conjecture 1, hold in this example. The atoms of $e,42131]_{R}$ are $1,2,4$ (the left descents), the atoms of $1,42131]_{R}$ are $41,21$ and the atom of $4231,42131]_{R}$ is $42131$, which illustrates the statement of Conjecture 3.
1.2. Conjecture 1 is true for finite Weyl groups
As stated above, Conjecture 1 is a generalization of a conjecture formulated by Ressayre in the case of finite Weyl groups, which the first author was made aware of in 2014. Let $W$ be a finite Weyl group with longest element $w_{\circ}$ and let $u_{1},u_{2},u_{3}\in W$, Ressayre conjectured the following statement:
$T(w_{\circ})=T(u_{1})\sqcup T(u_{2})\sqcup T(u_{3})\implies\operatorname{rk}(W)=|S|=d_{R}(u_{1})+d_{R}(u_{2})+d_{R}(u_{3})\qquad(\triangle).$
In other words, if $\{u_{1},u_{2},u_{3}\}$ is a $3$-partition of $w_{\circ}$ then $|S|=d_{R}(u_{1})+d_{R}(u_{2})+d_{R}(u_{3})$.
In 2023, Francone and Ressayre [10, Theorem 1] showed that the structure coefficients of the Belkale-Kumar product are either 0 or $1$ for any finite Weyl group, answering by the positive [5, Claim 1]. Their theorem implies $(\triangle)$ as a corollary [10, Corollary 5]. Their proof relies on an algebraic geometry argument, based on the fact that complete flag varieties are simply connected, to reduce the problem to a case-by-case computation on crystallographic root systems spanning over twenty pages. It turns out that Conjecture 1 is equivalent to statement $(\triangle)$, see Proposition 2.7. So Conjecture 1 is true for finite Weyl groups as a consequence of [10, Theorem 1].
Unfortunately, this equivalent statement $(\Delta)$ is not valid for infinite Coxeter systems (since the longest element $w_{\circ}$ is involved), whereas Conjecture 1 seems to be. So direct or undirected proofs of $(\Delta)$ won’t likely be generalizable to arbitrary Coxeter system. But still a direct proof of statement $(\Delta)$ for finite Coxeter groups could be interesting to better understand Conjecture 1.
In type $A$, a proof [^2] The authors of [4] states that their result can be extended to type $B$. of $(\Delta)$, involving solely the structure of the corresponding crystallographic root system, is given thanks to [4, Proposition 8.1]. The idea is as follows. For $w\in W$, denote by $\Phi^{R}(w)=\{-w(\alpha_{s})\mid s\in D_{R}(w)\}$, the so called set of descent-roots. In [4, Proposition 8.1], the authors show that if $\{u_{1},u_{2},u_{3}\}$ are a $3$-partition of $w_{\circ}$, then $\Phi^{R}(u_{1})\sqcup\Phi^{R}(u_{2})\sqcup\Phi^{R}(u_{3})$ is a basis of the vector space spanned by $\Phi$, which is of dimension $|S|$. Hence
$|S|=|\Phi^{R}(u_{1})|+|\Phi^{R}(u_{2})|+|\Phi^{R}(u_{3})|=d_{R}(u_{1})+d_{R}(u_{2})+d_{R}(u_{3}).$
Unfortunately, we do not see how to adapt this argument in the case of Conjecture 1. Indeed, in Example 1.4 (type $A_{3}$), the element $w=\tau_{1}\tau_{3}\tau_{2}\tau_{1}\tau_{3}$ has $u=\tau_{1}\tau_{2}$ and $v=\tau_{3}\tau_{2}\tau_{1}$ as a bipartition. Consider the crystallographic root system of $A_{3}$ with positive roots $e_{j}-e_{i}$ with $1\leq i<j\leq 4$. The unique inversion-descent of $u$ is $e_{3}-e_{1}$, the unique one for $v$ is $e_{4}-e_{1}$ - and those form a base of a subspace $V_{1}$. However, the inversion descents of $w$ are $e_{3}-e_{1}$ and $e_{4}-e_{2}$, which form a basis of a subspace $V_{2}$, distincts of $V_{1}$. If $V_{1}$ would have been equal to $V_{2}$, we could have concluded as in [4, Proposition 8.1]. This phenomenon is not an exception and happens in most cases of bipartitions. It would be interesting to know if a statement similar to [4, Proposition 8.1] holds for bipartitions of elements of $w$. The proof of [4, Proposition 8.1] is not elementary and is a consequence of a long and technical analysis of some properties of crystallographic root systems of type $A$.
@@ Acknowledgements
Part of this research was done at the Laboratoire d’algèbre, de combinatoire et d’informatique mathématique (LACIM) at UQAM during the sabbatical of the second author. Therefore, she warmly thanks the IRL CRM-CNRS of Montréal and the INSMI for giving her this opportunity. The first author warmly thanks Riccardo Biagioli, Matthew Dyer, Philippe Nadeau, Piotr Przytycki, Nicolas Ressayre and Paolo Sentinelli for instructive discussions.
@@ 2. Partitions of elements in a Coxeter system
In the weak order $(W,\leq_{R})$, it is well-known that, for $u,w\in W$, we have $u\leq_{R}w$ if and only if $T(u)\subseteq T(w)$. Björner showed that $(W,\leq_{R})$ is a complete meet-semilattice. For $X\subseteq W$, we denote the meet of $X$ by $\bigwedge_{R}X$ and, if it exists, the join of $X$ by $\bigvee_{R}X$.
@@ 2.1. First properties of partitions of elements of $W$
(1) The statement is a direct consequences of the definition of partitions of $w$. (2) The statement is a consequence of the definition of partitions of $w$ and of the fact that $|T(g)|=\ell(g)$ for any $g\in W$. (3) The statement is a consequence of the definition of partitions of $w$ and of the fact that $u\leq_{R}w$ if and only if $T(u)\subseteq T(w)$. (4) Assume by contradiction that $\wedge_{R}\mathcal{P}\not=e$. Then there is $s\in S$ such that $s\leq_{R}u$ for all $u\in\mathcal{P}$. So
@@
Recall that $(W,S)$ is a universal Coxeter system if its Coxeter graph is a complete graph whose edges are all indexed by $\infty$. The following proposition states that universal Coxeter systems satisfy Conjecture 1, which is obviously true if there are only partition-irreducible elements in $W$.